OpenStudy (anonymous):

find the rate of change for 1995 to 2000 given these ordered pairs in the data set" 1995,800) and (2000,1200). a.5 b.400 c.80 d.1200

OpenStudy (anonymous):

80

OpenStudy (anonymous):

(1200 - 800)/(2000-1995) = 400/5 = 80

OpenStudy (saifoo.khan):

Eighty,